OpenStudy (anonymous):

The graph of cosy=x-4y, what is the range of slopes of its tangent lines?

OpenStudy (aum):

\[ \cos y=x-4y \\ \text{Differentiate implicitly:} \\ -\sin(y) y' = 1 - 4y' \\ 4y' - \sin(y)y' = 1 \\ y'(4 - \sin(y)) = 1 \\ y' = \frac{1}{4 - \sin(y)} \\ \text{Range of }\sin(y) \text{ is: }[-1, 1] \\ \text{Therefore, range of }y' \text{ is: }\frac{1}{4 + 1} \text{ to } \frac{1}{4 - 1} = \large \left [\frac 15, \frac 13 \right ]. \\ \]